Solve l 2-3x l <=4...Solve absolute value sign 2-3x absolute value sign is less than or equal to 4.
Answer by stanbon(75887)

Solve l 2-3x l <=4
------
-4 <= 2-3x <= 4
------
Subtract 2 along the line to get:
-6 <= -3x <= 2
----
Divide thru by -3.
Remember: the inequality reverses:
-2/3 <= x <= 2
===================
Cheers,
Stan H.